OpenStudy (anonymous):

What is a quartic function with only the two real zeros given? x = -4 and x = -1

OpenStudy (anonymous):

Just punch in the values given for x into each function, and which ever one returns 0 for both values of x satisfies the condition. so for example: (-4)^4+5(-4)^3+5(-4)^2+5(-4)+4= 256-320+80-20+4=0 And (-1)^4+5(-1)^3+5(-1)^2+5(-1)+4= 1-5+5-5+4= 0